paul can walk 2 1/2 miles in 1 hour. How far can he walk in 1 3/4 hours?
Found 2 solutions by rfer, edjones:
Answer by rfer(16322)

2.5/1=x/1.75
x=4.4 miles

Answer by edjones(8007)

x/(7/4) = (5/2)/1
x= 7/4 * 5/2
= 35/8 mi
.
Ed